1. An unknown resistor (R3) is connected in series with the two other resistors (R1= 20 ohm and R2= 25 ohm) across a 15 V battery in the order R1arrowR2arrowR3. An ammeter, which is placed in between the battery and R1, reads a current of 200 mA.

a. what is the total resistance of the the circuit?

b. what is the resistance of the inknown resistor, R3?

c. what is the total current?

d. what is the current through each resistor?

e. what is the voltage drop across each resistor?

f.   Draw and label the circuit

2. Describe the anatomy of movement of the wave of cell depolarization throughout the heart during a cardiac cycle.

3.what is the cardiac output for a person whose heart beats 70 beats/min with a stroke volume of 70mL?

(Bioengineering)
Answer #1

1 a.

V= IR

R= V/I = 15/.2 = 75 ohm

1 b.

R_{1}+R_{2} + R_{3}= R

20+25 + R_{3}= 75

R_{3}= 30 ohm

1 c. Total current = 200mA

1 d .Current through each resistor = 200mA (as all are in series)

1 e. Votage drop across each resistor

V_{1} = I R_{1}

= .2 * 20

=4V

V_{2} = I R_{2} = .2 * 25 = 5V

V_{3} = I R_{3} = .2 * 30 = 6V

Circuit

Rest both are not Physics question